# KiDS-Legacy: angular galaxy clustering from deep surveys with complex selection effects

### v1

@@ 7 Discussions and conclusions
This work aims to correct the bias in the galaxy 2PCF from deep surveys due to complicated selection effects. We show that for deep surveys such as KiDS, the selection effect is very pronounced and can lead to a bias in the 2PCF about 10 times the signal, especially on large scales. Therefore, a critical step is to correct for the variable depth caused by complex selection effects. We introduce the SOM+HC method proposed by Johnston et al. (2021) on the KiDS-Legacy sample and show that it can correct the bias for faint galaxy samples.
@@
SOM+HC is a machine-learning method for recovering organised randoms that have the same selection pattern as the galaxy sample. The method uses a combination of self-organising maps and hierarchical clusters that group galaxies in systematics space and redistribute them across the survey footprint. Compared to other methods used to mitigate selection bias in the 2PCF, SOM+HC has the advantage of being model-independent. As an unsupervised machine learning algorithm, it does not need to parameterise selection functions or biases in statistics but directly captures arbitrary selection-induced patterns in the systematics space. Therefore, it can recover complex selection functions and their correlations without making a priori assumptions about them. In addition, the SOM+HC method is purely data-driven, meaning that it can be trained on the real data itself without relying on mock data or any external information.
@@
Johnston et al. (2021) proved that SOM+HC can correct the slight bias in the bright sample, while this work further validates the method on a faint galaxy sample with more complex selection effects and a more significant bias in the 2PCF. The validations are performed on mock galaxies from GLASS with both toy selections and data-driven selections. The toy test is a direct demonstration of how SOM+HC can recover the selection functions and the unbiased 2PCF. The data-driven test demonstrates the performance of SOM+HC on deep surveys such as KiDS and determines the optimal setup for realistic 2PCF measurements. From Figure 14, we note that the bias in the “UR” 2PCF is scale-dependent. This scale dependency may be because different scales are dominated by selection effects of different systematics, but it is hard to quantitatively isolate them in different scales. OR is able to correct such complicated selection effects. From this test, we see that the selection effects in the real KiDS-Legacy galaxy sample bias the 2PCF by $\chi^{2}\sim 8500$ and that the SOM+HC with $NC=600$ can correct this bias down to a level of $\chi^{2}\sim 0.3$ given a degree of freedom of 10, resulting in the reduction of bias on the $(\Omega_{\rm m},b)$ from a level of $27\sigma$ to $0.3\sigma$. Notably, this setup can recover an unbiased 2PCF even with an enhanced selection, correcting a 2PCF bias of $\chi^{2}\sim 20000$ to $\chi^{2}\sim 0.4$.
@@
Several setups of the SOM+HC algorithm can affect its performance, so the optimal setup varies from survey to survey. For example, the number of galaxies (data volume) is important for the SOM+HC setup. A catalogue with a larger number of galaxies is less sensitive to Poisson noise and has a more obvious variable depth, making the SOM+HC more efficient in recovering an accurate OR. On the other hand, smaller galaxy catalogues require lower $NC$ and Nside to reduce cosmological contamination in the OR, but this will limit the spatial resolution of the OR. Furthermore, SOM+HC performed on a dataset with weaker selection effects than the LSS fluctuation (such as the KiDS bright sample) is more prone to overcorrection. Therefore, SOM+HC should perform better on larger, fainter galaxy samples. For KiDS-Legacy, we have shown that $NC=600$ gives a parameter constraint accuracy of $\sim 0.3\sigma$ and performance is found to be stable around this choice. In this work we have also set the pixel size to 1.4 arcmin, equivalent to Nside=2048, to avoid unintentionally masking pixels within the footprint, while maintaining the best performance of the OR. We expect smaller pixels to give a more accurate OR for future surveys with higher galaxy number densities.
@@
Using the validated SOM+HC method, we perform a preliminary blinded 2PCF measurement from the KiDS-Legacy galaxy catalogue. We find that the UR 2PCF is significantly higher than the OR 2PCF by an order of magnitude. Furthermore, the corrected 2PCF is robust to the choice of $NC$ around 600. We will apply the SOM+HC method to the photometric galaxy clustering measurement for the KiDS-Legacy 6$\times$2pt and the bright galaxy clustering measurement for the KiDS-Legacy 3$\times$2pt measurements. More detailed discussions of the methodology (e.g. tomographic galaxy clustering) are given in the forthcoming papers.
@@
We moreover note that the tests carried out in this work are focused on linear scales. We have found that the performance of the method is worse on small scales. This is because the variance is very small on small scales, so the bias tends to be significant compared to it. Therefore, we need to take extra care at these scales. For example, future work could combine different algorithms to correct for variable depth at different scales.
@@
Combinations of multiple 2PCFs, namely the $N\times 2$pt measurements, will form the main analysis components of the next generation of LSS surveys, including LSST and Euclid. Galaxy 2PCF will be a critical part of these measurements. For these deep surveys, selection effects will also lead to variable depths. For example, the “rolling cadence” survey strategy of LSST (Bianco et al., 2021) will introduce stripe-like non-uniformity. Based on Hang et al. (2024), this will cause a significant bias up to an order of magnitude for the LSST Y3 galaxy clustering measurement. The Euclid survey, on the other hand, will combine ground-based multiband photometry to estimate photo-$z$ (Euclid Collaboration et al., 2020), so any photo-$z$-based selection will introduce variable depth. With large sky coverage and depth, the measurement precision of these surveys will be greatly improved, requiring cleaner and more reliable bias correction. As discussed above, SOM+HC is more effective with larger galaxy samples. Therefore, we expect that as the data volumes of the next generation of surveys increase, SOM+HC will become more powerful in recovering selection-induced clustering. In addition, one can use smaller pixels (or higher Nside) for such galaxy catalogues, and obtain the OR weights at higher angular resolutions. To this end, we publish the code used in this work as the tiaogeng package 151515 https://github.com/yanzastro/tiaogeng/ for future implementation in pipelines for next-generation surveys (e.g. TXPipe [^16] https://txpipe.readthedocs.io/en/latest/). A combination with other methods, such as template-based correction methods, could be even more effective in mitigating the complex selection effects for future deep surveys.
@@ Acknowledgements.
The data in this paper is analysed with open-source python packages numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), astropy (Astropy Collaboration et al., 2018), matplotlib (Hunter, 2007), GLASS(Tessore et al., 2023), somoclu(Wittek et al., 2017), healpy (Zonca et al., 2019), NaMaster (Alonso et al., 2019), CCL (Chisari et al., 2019), emcee (Foreman-Mackey et al., 2013), and GetDist (Lewis, 2019). We also use WebPlotDigitizer (Rohatgi, 2021) to digitize some external data from plots in the literature.
@@ Appendix A Covariance matrix comparison
In this appendix section, we compare the covariance matrix from mock realisations and the theoretical covariance computed by the OneCovariance code. The mock covariance is computed based on Eq. (17). To ensure consistency, we configure the input file of the OneCovariance code so that the cosmology, footprint, redshift distribution and galaxy number density match those of the GLASS mock. In addition, since the GLASS package only generates a Gaussian field, we assume that the theoretical covariance matrix contains only the Gaussian covariance plus the super-sample covariance (Takada & Hu 2013) to account for the covariance of modes larger than the observed field.
In this section we compare only the covariance of the 2PCFs from the data-driven test with $\{NC_{\mathrm{KiDS}}=600,NC_{\mathrm{rec}}=600\}$. The three panels in Figure 20 show the square root of the main diagonal, the 5th and the 10th diagonal above the mean diagonal. Note that the covariance of the uniform random case is significantly higher than that of the “no selection” case, implying that variable depth contamination also introduces additional covariance into the data. The covariance for the “No selection” and “True OR” cases matches the theoretical covariance given by OneCovariance as expected. The covariance of “Recovered OR” agrees with that of the unbiased cases, so we conclude that the SOM+HC method can also recover accurate covariance in the correlation function in the linear regime.
Figure 20: The comparison of the covariance matrices. Panels from top to bottom show the square root of the main covariance diagonal, the 5th and 10th diagonals above the main diagonal, as shown by the black grids in the top right theoretical covariance matrices in each panel. Coloured curves show the covariance terms of the four 2PCFs defined in Table 2 from the data-driven test with $\{NC_{\mathrm{KiDS}}=NC_{\mathrm{rec}}=600\}$; the dotted lines are calculated from the OneCovariance code. The lower x-ticks are the row indices, while the upper x-ticks on the top panel are the corresponding angular scale for the diagonal term.
@@ Appendix B Correcting selection effects for angular power spectra with SOM+HC
Figure 21: Upper panel: Pseudo-$C_{\ell}$ measured from GLASS mock samples. The data points of each series are the average PCL from 40 realisations; the error bars are the square root of the covariance diagonal given by NaMaster. The $\chi^{2}_{\mathrm{d}}$ values, which describe the difference between each data set and the “no selection” PCL, are calculated similarly to Eq. (18). Lower panel: The relative difference of each case with respect to the “no selection” case. The shaded regions are the $\ell$ modes corresponding to a physical scale smaller than 8 $h^{-1}$pc estimated at the mean redshift.
@@
Another widely used 2-point statistic is the angular power spectrum, which is defined as the correlation of galaxy overdensity in harmonic space. In general, angular power spectra have weak correlations between $\ell$ modes, allowing us to study different angular scales independently.
@@
The other estimator is the pseudo-$C_{\ell}$ (PCL, Wandelt et al. 2001; Alonso et al. 2018), i.e. one calculates the “coupled” $C_{\ell}$ from the weighted sky map, and then “decouple” it with the mode coupling matrix of the weight and bin it. The weight can be a binary mask specifying the footprint of the survey (so that a source gets a weight of 1 if it is in the footprint, otherwise 0); or a weight for each source to suppress errors (like the lensing weight) or to correct for selection effects (like the organised random weight in this paper). In this appendix, we briefly discuss the usage and performance of SOM+HC on the measurement of PCL $C_{\ell}$ from the same GLASS mock samples. Detailed discussions will be given in a companion paper presenting the methodology of the KiDS-Legacy 6$\times$2pt cosmology.
The galaxy PCL is based on pixelised galaxy overdensity maps. On a weighted sky, the galaxy overdensity in the $p$-th pixel is given by (Nicola et al. 2020):
@@
where $M_{\ell\ell^{\prime}}(w^{a},w^{b})$ is the mode-mixing matrix determined by the weight maps $(w^{a},w^{b})$ of the two fields. Therefore, the directly measured $\tilde{C}^{ab}_{\ell}$ is called “coupled PCL”. An unbiased angular power spectrum is estimated by decoupling it via the inverse of the mode-coupling matrix. Assuming Poisson-distributed objects, the OR weight is proportional to the inverse variance of the field, we use OR weight to calculate the mode-coupling matrix. In practice, one also wants to bin the PCL into $\ell$ bins. For technical details, we refer the readers to Alonso et al. (2019). The calculations are implemented in the NaMaster package (Alonso et al. 2019) which we will use in this section to measure the PCL.
@@
The Gaussian covariance matrix depends on unbiased estimations of the angular power spectra. In this section, we only measure angular power spectra from GLASS mock catalogue, so we take the input angular power spectra for GLASS to estimate the theoretical Gaussian covariance matrix. [^17] In real measurement when the true angular power spectra are unknown, an iterative estimation is usually used to evaluate the covariance (Eifler et al. 2009). That is, one uses the theoretical angular power spectra calculated from reasonable cosmological parameters to calculate the covariance and constrain the parameters with the corresponding likelihood. The covariance matrix is updated with the best-fit parameter. This process is performed iteratively until the best-fit parameter converges. The mode-coupling induced by the OR weights is also taken care of by the NaMaster package. The non-Gaussian term includes a connected covariance matrix which is dominated by the galaxy trispectra (Krause & Eifler 2017) which only affects the small scales, so we neglect it in this section. Another non-Gaussian term is the super-sample covariance (SSC; Takada & Hu 2013) which accounts for the correlated modes that are larger than the survey footprint. We calculate this term using the OneCovariance package, but we note that this term is negligible for KiDS-Legacy.
@@
Therefore, for PCL, the variable depth affects the mapping process, the mode-mixing matrix, the shot noise and the covariance matrix. In this section, we perform the data-driven test for PCL with the same 40 GLASS mock samples with interpolated systematics. Here we show the case $NC_{\mathrm{KiDS}}=NC_{\mathrm{rec}}=600$. That is, the mock sample is selected with the data-driven OR weight with $NC_{\mathrm{KiDS}}=600$ and the recovered OR weight is from 600 HCs trained on the selected mock sample. We first generate the galaxy fluctuation map according to Eq. (24) and measure the four PCLs defined similarly to Table 2: the “UR” case is a depleted galaxy sample for which the mode-mixing matrix is calculated from the footprint; the “No selection” case is the unselected galaxy sample for which the mode-mixing matrix is calculated from the footprint, etc.
@@
The PCL measure is shown in Figure 21. Data points in the top panel show the averaged PCL across realisations. Error bars are the standard deviation of each $C_{\ell}$ calculated from the covariance matrix given by NaMaster. The bottom panel shows the relative difference between each $C_{\ell}$ measurement and the “no selection” case. The shaded region corresponds to physical scales smaller than $8h^{-1}\mathrm{Mpc}$ estimated at the mean redshift. We ignore these scales in this analysis. To assess the consistency between the OR-corrected PCL and the “no selection” PCL, we calculate the $\chi_{\mathrm{d}}$ (defined similarly to Eq. (18) but using $C_{\ell}$ instead of $w(\theta)$) between them. The values are quoted in the upper panel of Figure 21. The uniform random again gives a very large bias ($\chi_{\mathrm{d}}=3230$) in the PCL and the true OR completely corrects the bias as expected. The OR recovered by SOM+HC gives a residual of $\chi_{\mathrm{d}}=0.42$.
@@
In the same way as for the 2PCF, we run an MCMC on the measured PCL. The Gaussian likelihood is defined with the Gaussian covariance given by NaMaster plus the SSC given by OneCovariance. We calculate the shifted posterior with respect to the best-fit values constrained from the “No selection” case. The contours of the 68% credible level are shown in Figure 22. We calculate the parameter constraint bias following the 2PCF test and get $\Delta\Omega_{\mathrm{m}}[\sigma]=0.21$, $\Delta b[\sigma]=-0.17$, $\chi^{2}_{\Delta\Omega_{\mathrm{m}},\Delta b}=0.07$ and for the recovered OR weight case. For the true OR case, the constraining bias is at the $0.001\sigma$ level. We note that both posteriors have the same shape, indicating a close degeneracy between the galaxy bias and $\Omega_{\mathrm{m}}$.

### v2

@@ 7 Discussions and conclusions
The aim of this work was to correct the bias in the galaxy 2PCF from deep surveys due to complicated selection effects. We showed that for deep surveys such as KiDS, the selection effect is very pronounced and can lead to a bias in the 2PCF about ten times the signal, especially on large scales. Therefore, a critical step is to correct for the variable depth caused by complex selection effects. We introduce the SOM+HC method proposed by Johnston et al. (2021) on the KiDS-Legacy sample and show that it can correct the bias for faint galaxy samples.
@@
SOM+HC is a machine-learning method for recovering organised randoms that have the same selection pattern as the galaxy sample. The method uses a combination of self-organising maps and hierarchical clusters that group galaxies in systematics space and redistributes them across the survey footprint. Compared to other methods used to mitigate selection bias in the 2PCF, SOM+HC has the advantage of being model-independent. As an unsupervised machine learning algorithm, it does not need to parametrise selection functions or biases in statistics, but directly captures arbitrary selection-induced patterns in the systematics space. Therefore, it can recover complex selection functions and their correlations without making a priori assumptions about them. In addition, the SOM+HC method is purely data-driven, meaning that it can be trained on the real data itself without relying on mock data or any external information.
@@
Johnston et al. (2021) proved that SOM+HC can correct the slight bias in the bright sample, while this work further validates the method on a faint galaxy sample with more complex selection effects and a more significant bias in the 2PCF. The validations are performed on mock galaxies from GLASS with both toy selections and data-driven selections. The toy test is a direct demonstration of how SOM+HC can recover the selection functions and the unbiased 2PCF. The data-driven test demonstrates the performance of SOM+HC on deep surveys such as KiDS and determines the optimal set-up for realistic 2PCF measurements. From Fig. 14, we note that the bias in the UR 2PCF is scale-dependent. This scale dependency may arise because different scales are dominated by selection effects of different systematics, but it is hard to quantitatively isolate them in different scales. OR is able to correct such complicated selection effects. From this test, we see that the selection effects in the real KiDS-Legacy galaxy sample bias the 2PCF by $\chi^{2}\sim 8500$ and that the SOM+HC with $NC=600$ can correct this bias down to a level of $\chi^{2}\sim 0.3$ given a degree of freedom of 10, resulting in the reduction of bias on the $(\Omega_{\rm m},b)$ from a level of $27\sigma$ to $0.3\sigma$. Notably, this set-up can recover an unbiased 2PCF even with an enhanced selection, correcting a 2PCF bias of $\chi^{2}\sim 20000$ to $\chi^{2}\sim 0.4$.
@@
Several set-ups of the SOM+HC algorithm can affect its performance, so the optimal set-up varies from survey to survey. For example, the number of galaxies (data volume) is important for the SOM+HC set-up. A catalogue with a larger number of galaxies is less sensitive to Poisson noise and has a more obvious variable depth, making the SOM+HC more efficient in recovering an accurate OR. On the other hand, smaller galaxy catalogues require lower $NC$ and Nside to reduce cosmological contamination in the OR, but this will limit the spatial resolution of the OR. Furthermore, data-driven regression methods, including SOM+HC, aim to nullify modes of variance in the training data, which carries the risk of over-correction. A potential solution is to train the model on data from a disjoint sky region and run it on the region of interest. This approach requires ensuring that the training region is both representative and uncorrelated with the testing region (the region of interest), making it practically challenging. For datasets with weaker selection effects than the LSS fluctuation (such as the KiDS bright sample), SOM+HC is more prone to over-correction. Therefore, SOM+HC should perform better on larger, fainter galaxy samples. For KiDS-Legacy, we show that $NC=600$ gives a parameter constraint accuracy of $\sim 0.3\sigma$ and performance is found to be stable around this choice. In this work we also set the pixel size to 1.4 arcmin, equivalent to Nside=2048, to avoid unintentionally masking pixels within the footprint, while maintaining the best performance of the OR. We expect smaller pixels to give a more accurate OR for future surveys with higher galaxy number densities.
@@
Using the validated SOM+HC method, we performed a preliminary blinded 2PCF measurement from the KiDS-Legacy galaxy catalogue. We find that the UR 2PCF is significantly higher than the OR 2PCF by an order of magnitude. Furthermore, the corrected 2PCF is robust to the choice of $NC$ around 600. We applied the SOM+HC method to the photometric galaxy clustering measurement for the KiDS-Legacy 6$\times$2pt and the bright galaxy clustering measurement for the KiDS-Legacy 3$\times$2pt measurements. More detailed discussions of the methodology (e.g. tomographic galaxy clustering) will be given in the forthcoming papers.
@@
We moreover note that the tests carried out in this work were focused on linear scales. The bottom panels in Fig. 14 show that the performance of the method is worse on small scales. This is because the variance is very small on small scales, so the bias tends to be significant compared to it. Therefore, we need to take extra care at these scales. For example, future work could combine different algorithms to correct for variable depth at different scales.
@@
Combinations of multiple 2PCFs, namely the $N\times 2$pt measurements, will form the main analysis components of the next generation of LSS surveys, including LSST and Euclid. Galaxy 2PCF will be a critical part of these measurements. For these deep surveys, selection effects will also lead to variable depths. For example, the rolling cadence survey strategy of LSST (Bianco et al., 2021) will introduce stripe-like non-uniformity. Based on Hang et al. (2024), this will cause a significant bias up to an order of magnitude for the LSST Y3 galaxy clustering measurement. The Euclid survey, on the other hand, will combine ground-based multi-band photometry to estimate photo-$z$ (Euclid Collaboration et al., 2020), so any photo-$z$-based selection will introduce variable depth. With large sky coverage and depth, the measurement precision of these surveys will be greatly improved, requiring cleaner and more reliable bias correction. As discussed above, SOM+HC is more effective with larger galaxy samples. Therefore, we expect that as the data volumes of the next generation of surveys increase, SOM+HC will become more powerful in recovering selection-induced clustering. In addition, one can use smaller pixels (or higher Nside) for such galaxy catalogues, and obtain the OR weights at higher angular resolutions. To this end, we published the code used in this work as the tiaogeng package191919 https://github.com/yanzastro/tiaogeng/ for future implementation in pipelines for next-generation surveys (e.g. TXPipe [^20] https://txpipe.readthedocs.io/en/latest/). A combination with other methods, such as template-based correction methods, could be even more effective in mitigating the complex selection effects for future deep surveys.
@@ Acknowledgements.
The data in this paper is analysed with open-source python packages numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), astropy (Astropy Collaboration et al., 2018), matplotlib (Hunter, 2007), GLASS(Tessore et al., 2023), somoclu(Wittek et al., 2017), healpy (Zonca et al., 2019), NaMaster (Alonso et al., 2019), CCL (Chisari et al., 2019), emcee (Foreman-Mackey et al., 2013), and GetDist (Lewis, 2019). We also use WebPlotDigitizer (Rohatgi, 2021) to digitise some external data from plots in the literature.
@@
Data availability: Our software is open-source for future usage.
@@ Appendix A Covariance matrix comparison
In this Appendix we compare the covariance matrix from mock realisations and the theoretical covariance computed by the OneCovariance code. The mock covariance is computed based on Eq. (17). To ensure consistency, we configure the input file of the OneCovariance code so that the cosmology, footprint, redshift distribution and galaxy number density match those of the GLASS mock. In addition, since the GLASS package only generates a Gaussian field, we assume that the theoretical covariance matrix contains only the Gaussian covariance plus the super-sample covariance (Takada & Hu 2013) to account for the covariance of modes larger than the observed field.
In this section we compare only the covariance of the 2PCFs from the data-driven test with $\{NC_{\mathrm{KiDS}}=600,NC_{\mathrm{rec}}=600\}$. The three panels in Fig. 20 show the square root of the main diagonal, the 5th and the 10th diagonal above the mean diagonal. We note that the covariance of the uniform random case is significantly higher than that of the “no selection” case, implying that variable depth contamination also introduces additional covariance into the data. The covariance for the “No selection” and “True OR” cases matches the theoretical covariance given by OneCovariance as expected. The covariance of “Recovered OR” agrees with that of the unbiased cases, so we conclude that the SOM+HC method can also recover accurate covariance in the correlation function in the linear regime.
Figure 20: Comparison of the covariance matrices. Panels from top to bottom show the square root of the main covariance diagonal, the 5th and 10th diagonals above the main diagonal, as shown by the black grids in the top right theoretical covariance matrices in each panel. The coloured curves show the covariance terms of the four 2PCFs defined in Table 16 from the data-driven test with $\{NC_{\mathrm{KiDS}}=NC_{\mathrm{rec}}=600\}$; the dotted lines are calculated from the OneCovariance code. The lower x-ticks are the row indices, while the upper x-ticks on the top panel are the corresponding angular scale for the diagonal term.
@@ Appendix B Correcting selection effects for angular power spectra with SOM+HC
Figure 21: Upper panel: Pseudo-$C_{\ell}$ measured from GLASS mock samples. The data points of each series are the average PCL from 40 realisations; the error bars are the square root of the covariance diagonal given by NaMaster. The $\chi^{2}_{\mathrm{d}}$ values, which describe the difference between each dataset and the No selection PCL, are calculated similarly to Eq. (18). Lower panel: Relative difference of each case with respect to the no selection case. The shaded regions are the $\ell$ modes corresponding to a physical scale smaller than 8 $h^{-1}$pc estimated at the mean redshift.
@@
Another widely used two-point statistic is the angular power spectrum, which is defined as the correlation of galaxy over-density in harmonic space. In general, angular power spectra have weak correlations between $\ell$ modes, allowing us to study different angular scales independently.
@@
The other estimator is the pseudo-$C_{\ell}$ (PCL, Wandelt et al. 2001; Alonso et al. 2018) , where one calculates the coupled $C_{\ell}$ from the weighted sky map, then decouples it with the mode coupling matrix of the weight, and bins it. The weight can be a binary mask specifying the footprint of the survey (so that a source gets a weight of 1 if it is in the footprint, otherwise 0); or a weight for each source to suppress errors (like the lensing weight) or to correct for selection effects (like the organised random weight in this paper). In this appendix we briefly discuss the usage and performance of SOM+HC on the measurement of PCL $C_{\ell}$ from the same GLASS mock samples. Detailed discussions will be given in a companion paper presenting the methodology of the KiDS-Legacy 6$\times$2pt cosmology.
The galaxy PCL is based on pixelised galaxy over-density maps. On a weighted sky, the galaxy over-density in the $p$-th pixel is given by (Nicola et al. 2020):
@@
where $M_{\ell\ell^{\prime}}(w^{a},w^{b})$ is the mode-mixing matrix determined by the weight maps $(w^{a},w^{b})$ of the two fields. Therefore, the directly measured $\tilde{C}^{ab}_{\ell}$ is called “coupled PCL”. An unbiased angular power spectrum is estimated by decoupling it via the inverse of the mode-coupling matrix. Assuming Poisson-distributed objects, the OR weight is proportional to the inverse variance of the field, we use OR weight to calculate the mode-coupling matrix. In practice, one also wants to bin the PCL into $\ell$ bins. For technical details, we refer to Alonso et al. (2019). The calculations are implemented in the NaMaster package (Alonso et al. 2019) which we use in this section to measure the PCL.
@@
The Gaussian covariance matrix depends on unbiased estimations of the angular power spectra. In this section we only measure angular power spectra from GLASS mock catalogue, so we take the input angular power spectra for GLASS to estimate the theoretical Gaussian covariance matrix. [^21] In real measurement when the true angular power spectra are unknown, an iterative estimation is usually used to evaluate the covariance (Eifler et al. 2009). That is, one uses the theoretical angular power spectra calculated from reasonable cosmological parameters to calculate the covariance and constrain the parameters with the corresponding likelihood. The covariance matrix is updated with the best-fit parameter. This process is performed iteratively until the best-fit parameter converges. The mode-coupling induced by the OR weights is also taken care of by the NaMaster package. The non-Gaussian term includes a connected covariance matrix which is dominated by the galaxy trispectra (Krause & Eifler 2017) which only affects the small scales, so we neglect it in this section. Another non-Gaussian term is the super-sample covariance (SSC; Takada & Hu 2013) which accounts for the correlated modes that are larger than the survey footprint. We calculate this term using the OneCovariance package, but we note that this term is negligible for KiDS-Legacy.
@@
Therefore, for PCL, the variable depth affects the mapping process, the mode-mixing matrix, the shot noise and the covariance matrix. In this section we perform the data-driven test for PCL with the same 40 GLASS mock samples with interpolated systematics. Here we show the case $NC_{\mathrm{KiDS}}=NC_{\mathrm{rec}}=600$. That is, the mock sample is selected with the data-driven OR weight with $NC_{\mathrm{KiDS}}=600$ and the recovered OR weight is from 600 HCs trained on the selected mock sample. We first generate the galaxy fluctuation map according to Eq. (24) and measure the four PCLs defined similarly to Table 16: the UR case is a depleted galaxy sample for which the mode-mixing matrix is calculated from the footprint; the ‘No selection’ case is the unselected galaxy sample for which the mode-mixing matrix is calculated from the footprint.
@@
The PCL measure is shown in Fig. 21. Data points in the top panel show the averaged PCL across realisations. Error bars are the standard deviation of each $C_{\ell}$ calculated from the covariance matrix given by NaMaster. The bottom panel shows the relative difference between each $C_{\ell}$ measurement and the “no selection” case. The shaded region corresponds to physical scales smaller than $8h^{-1}\mathrm{Mpc}$ estimated at the mean redshift. We ignore these scales in this analysis. To assess the consistency between the OR-corrected PCL and the “no selection” PCL, we calculate the $\chi_{\mathrm{d}}$ (defined similarly to Eq. (18) but using $C_{\ell}$ instead of $w(\theta)$) between them. The values are quoted in the upper panel of Fig. 21. The uniform random again gives a very large bias ($\chi_{\mathrm{d}}=3230$) in the PCL and the true OR completely corrects the bias as expected. The OR recovered by SOM+HC gives a residual of $\chi_{\mathrm{d}}=0.42$.
@@
In the same way as for the 2PCF, we run an MCMC on the measured PCL. The Gaussian likelihood is defined with the Gaussian covariance given by NaMaster plus the SSC given by OneCovariance. We calculate the shifted posterior with respect to the best-fit values constrained from the “No selection” case. The contours of the 68% credible level are shown in Fig. 22. We calculate the parameter constraint bias following the 2PCF test and get $\Delta\Omega_{\mathrm{m}}[\sigma]=0.21$, $\Delta b[\sigma]=-0.17$, $\chi^{2}_{\Delta\Omega_{\mathrm{m}},\Delta b}=0.07$ and for the recovered OR weight case. For the true OR case, the constraining bias is at the $0.001\sigma$ level. We note that both posteriors have the same shape, indicating a close degeneracy between the galaxy bias and $\Omega_{\mathrm{m}}$.